Calculate the number of grams-formula weights in 500 g of methyl alcohol, CH3OH


1
Expert's answer
2021-02-28T06:15:37-0500

n = m/M

M (CH3OH) = 32.04 g/mol

n (CH3OH) = 500/32.04 = 15.61 mol